undeadknight26 (undeadknight26):

Am I right? What are the explicit equation and domain for a geometric sequence with a first term of 4 and a second term of −12? an = 4(−3)n − 1; all integers where n ≥ 1<--- an = 4(−3)n − 1; all integers where n ≥ 0 an = 4(36)n − 1; all integers where n ≥ 1 an = 4(36)n − 1; all integers where n ≥ 0

undeadknight26 (undeadknight26):

@AverageJoe20 @ganeshie8 @JoannaBlackwelder @Lady.Liv1776 @narissa @Michele_Laino @TuringTest @thomaster @triciaal @Ineedhelpfast12

ganeshie8 (ganeshie8):

Correct. first term = 4 common ratio = -12/4 = -3 therefore the general term is given by \[a_n=4(-3)^{n-1}~;~n\ge 1\]

undeadknight26 (undeadknight26):

Can you check a few more sire?